Differential of log^n x 

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
f(x) = log^{n}(x)
f(x) = \frac{n.log^{n-1}(x)}{x}
t = log(x)
f(x) = t^{n}
f'(x) = n.t^{n-1}.\frac{dt}{dx}
\frac{dt}{dx} = \frac{1}{x}
f'(x) = n.t^{n-1}.\frac{1}{x}
f'(x) = n.(log(x))^{n-1}.\frac{1}{x}
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty